A random variable X is normally distributed with a mean of 2 and a standard deviation of 1.4.

Calculate the point c such that P ( X ≥ c ) = 0.5.
Answer #1

Solution:-

Given that,

mean = = 2

standard deviation = = 1.4

Using standard normal table,

P ( X ≥ c ) = 0.5

= 1 - P ( X < c ) = 0.5  

= P ( X < c ) = 1 - 0.5

=P ( X < c ) = 0.5

= P(Z < 0 ) = 0.5  

Using z-score formula,

x = z * +

x = 0 * 1.4+2

x = 2

answer =2